OpenStudy (anonymous):

The grocery store sells dates for $4.00 a pound and pomegranates for $2.75 a pound. Choose an equation in standard form for the weights of dates d and pomegranates p that a customer could buy with $12. A. 4p + 2.75d = 12 B. 4d = 2.75p + 12 C. 4d + 2.75p = 12 D. 4 + 2.75 = d @Mathematics HELP ME PLEASEE I THINK ITS C , What Do You Guys Think?

OpenStudy (anonymous):

anybodyy ?

OpenStudy (anonymous):

I'm just going to walk you through it $4 d (dates) $2.75 p (pomegranates) a certain amount of each will equal 12 that's multiplication, or 4d and 2.75p so your equation will be 4d+2.75p=12 because, if you add up the amount it cost for the dates (4d) and the amount it cost for the pomegranates (2.75p) it will equal $12 so the answer is: c. 4d + 2.75p = 12 (pay attention to which variable goes with what amount per pound)

OpenStudy (anonymous):

thanksss : )